OpenStudy (anonymous):

Choose the correct answer. Sean wants to make a mixture that is 50% lemon juice and 50% lime juice. How much 100% lemon juice should he add to a juice that is 30% lemon juice and 70% lime juice to make 7 gallons of the 50% lemon/50% lime juice mixture? A. 0.5 gallon B. 1 gallon C. 2 gallons D. 2.5 gallons Would this be 1 gallon..?

OpenStudy (whpalmer4):

He has two kinds of juice: let's call one \(s\) (for straight lemon juice) and the other \(m\) (for mixture of lemon and lime). We know that \[s+m = 7\]We also know that the amount of lemon juice in 1 unit of \(s\) is 1, or the lemon juice content of \(s\) is simply \(1s\). The amount of lemon juice in 1 unit of \(m\) is \(0.3m\). In the final mixture, we are going to have 0.5*7 = 3.5 units of lemon juice. \[1s + 0.3m = 3.5\] That gives you a system of two equations in two unknowns. Solve for \(s,m\) to get your answer.

OpenStudy (whpalmer4):

\(s\) will be the amount of 100% lemon juice added to the mixture to make the 50/50 lemon/lime juice mixture.

OpenStudy (whpalmer4):

You can verify that 1 gallon is not the correct answer: 1 gallon of lemon juice + 6 gallons of mixture gives you 1 gallon of lemon juice + 0.3*6 gallons = 1+1.8 = 2.8 gallons of lemon juice in 7 gallons of mixture. That isn't a 50/50 blend, clearly.

OpenStudy (praxer):

x= amount of pure lemon juice y= amount of gallons of 30/70 mixture 7= total amount of gallons of 50/50 mix x+y=7 Amount of lemon juice in p gallons of 100% lemon juice = 1x Amount of lemon juice in m gallons of 30% lemon juice = .3y Amount of lemon juice in 7 gallons of 50% lemon juice = 3.5 so x+.3y=3.5 Now solve: x+y=7 x=7-y substitute 7-y into second equation for x (7-y) +.3y=3.5 -.7y=-3.5 y=5 substitute 5 for y into the equation x=7-y x=7-5 x=2 gallons of 100% lemon juice I think so,

OpenStudy (praxer):

answer must be 2 gallons. If I am not wrong.

OpenStudy (anonymous):

I think that might be right

OpenStudy (whpalmer4):

Let's check: 2 gallons of 100% lemon juice added to 5 gallons of 30/70 mixture gives us 2 + 0.3*5 = 2+1.5 = 3.5 gallons of lemon juice and 0.7*5 = 3.5 gallons of lime juice, which is a 50/50 mixture, as desired.